OpenStudy (anonymous):

NEED TO CHECK SOME ANSWERS! √x+6=-2 Answer is x=-64 12√72+6√72= Answer is 18√72

OpenStudy (anonymous):

Checked.

OpenStudy (anonymous):

THey are right?

OpenStudy (anonymous):

The second is correct, but be careful about the first. Can you take the square of a negative number? x = 64 since the square root of 64 is either positive or negative 8.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

\[\sqrt{x+6}=-2??\]

OpenStudy (anonymous):

I was assuming it was \[\sqrt{x}+6 = -2\]

OpenStudy (anonymous):

yes both under the radical x+6=-2

OpenStudy (anonymous):

Problem √x+6 = -2 worked √x=-6-2, √=-8, (√x)^2 = -8^2, x=-64

OpenStudy (saifoo.khan):

Hi there.

OpenStudy (anonymous):

\[\sqrt{x+6}= -2\]\[x +6 =4\] x = -2